# Adaptive Plan-Execute Framework for Smart Contract Security Auditing

### v1

@@ 3.3.1 Execution and Reporting (A3-A5)
The final stage focuses on producing a clear, concise, and actionable audit report. This report consolidates the initial contract understanding ($s_{1}$), the audit plan ($s_{2}$), and the synthesized findings ($s_{4}$), including severity assessments and, where possible, actionable recommendations for mitigation. The structure and language of the report can be tailored to suit different audiences (e.g., developers requiring technical details, or management needing high-level summaries). Drawing insights from best practices in security reporting and relevant studies liu2024exploring; yin2024multitask, the report aims to cover key dimensions such as clear vulnerability descriptions, severity evaluations, precise localization of issues in the code, and concise summaries.
@@ 3.3.2 Task-Specific Review and Calibration
Given that LLMs can occasionally "hallucinate" or generate outputs that deviate from factual contract code bender2021dangers. Stage A3 is explicitly structured into two critical steps for each sub-task: (1) initial task-specific review and (2) rigorous calibration. This two-step process is designed to maximize accuracy and minimize false positives or missed vulnerabilities before findings are aggregated.
@@ 3.3.3 RAG-Enhanced Calibration for Improved Accuracy and Grounding
While pre-trained LLMs possess extensive general knowledge, their understanding of domain-specific knowledge can be limited or outdated. This can lead to gaps in analysis or less precise explanations, particularly for edge-case scenarios. To mitigate this, we enhance the calibration step with Retrieval-Augmented Generation (RAG) lewis2020retrieval, as illustrated in Figure 2. RAG augments the LLM’s internal knowledge by providing access to relevant excerpts from external, up-to-date knowledge sources during generation.
We have observed instances where an LLM might correctly identify a vulnerable code location but provide an incomplete or imprecise explanation, especially for complex or uncommon vulnerabilities. RAG addresses this by retrieving documents that offer detailed, authoritative explanations for similar issues. Within our Execute Phase, each calibration prompt $\mathcal{P}_{v}^{i}$ is augmented with information retrieved from a curated knowledge base. This knowledge base includes smart contract databases (e.g., SWC Registry, CWE, CVE), security guidelines (e.g., Ethereum Foundation, ConsenSys Diligence), seminal research papers, technical blogs detailing novel exploits, and community-vetted best practice repositories (see Appendix B for a list of sources). These resources are embedded and stored in a Vector Database (using models like Sentence-BERT reimers2019sentence) for efficient semantic retrieval.
@@ 3.4 LLM-Powered Audit Evaluator
Assessing the quality of the generation is an even more arduous task than the generation itself, and this issue has not been given adequate consideration. Few studies have pointed that LLM has been a good evaluator for natural language generation (NLG), and the existing approaches are not tailored to the specific needs of the audit process fu2023gptscore; liu2023g. Inspired by RAGAs es2024ragas, we designed and implemented an automated LLM-Powered Audit Evaluator. This evaluator systematically compare the findings generated by SmartAuditFlow against expert-annotated ground truth answers from the same expert audited-report. The primary objectives of this evaluator are to ensure fair and reproducible assessments, accelerate the evaluation cycle, and provide actionable insights for the continual improvement of the SmartAuditFlow framework. Our LLM-Powered Audit Evaluator operates in two main phases: (1) LLM-Driven Finding Comparison and (2) Quantitative Performance Scoring.
@@
  • Match Categorization: Based on the multi-dimensional assessment and the rubric, the LLM categorizes the relationship between the paired findings. Inspired by prior research wei2024measuring, we adopt a nuanced classification approach rather than a binary true/false distinction. The primary categories include Exact Match (all critical aspects align), Partial Match (core issue identified but differs on a specific dimension like severity or location precision), and Incorrect (the response misidentifies the core issue in the Standard Answer).
@@ 4.2.1 Dataset
  • Standard Vulnerability Set: This dataset contains smart contracts annotated with common vulnerability types, which serves as a benchmark for evaluating the detection of well-understood vulnerability types against established tools. We utilize the SmartBugs-curated dataset durieux2020empirical, which is widely adopted by both developers and researchers. The dataset includes 143 contracts, each annotated with common vulnerabilities based on the DASP classification, making it an excellent resource for benchmarking vulnerability detection tools.
  • Real-world Contracts Set: To assess performance on more complex, real-world scenarios, we use a curated subset of contracts from projects audited by Code4rena durieux2020empirical. This set consists of 72 projects, encompassing 6,454 contracts, among which 243 contracts have known issues, including 784 instances of high or medium-severity vulnerabilities. This dataset tests SmartAuditFlow’s capability to handle complex contract logic and diverse vulnerability manifestations.
  • CVE Set: We curated a set of Common Vulnerabilities and Exposures (CVEs) relevant to smart contracts. As of early 2025, there are 592 smart contract-related CVEs, with integer overflows being a predominant category.Followed by another co-worker’s PropertyGPT liu2024propertygpt, we selected 13 representative CVEs for detailed analysis, including three integer overflow cases, three access control vulnerabilities, and a diverse set of four other logic bugs.
@@ 4.3 Implementation Details
All experiments for SmartAuditFlow and static tools were conducted within a Docker container to ensure a consistent, isolated, and reproducible environment across all test runs. For evaluating the vulnerability detection capabilities involving LLMs, we adopted a single attempt approach (often denoted as pass@1) wei2024measuring. This reflects practical auditing scenarios where auditors typically seek an effective assessment in one primary attempt.
@@ 4.4.3 RQ3: Performance on Real-World Smart Contracts
Based on findings from previous experiments (RQ1, RQ2) and their prevalence, we selected a representative set of standalone LLMs for comparison: General-Purpose models: DeepSeek-v3, Gemini-2.0-flash, Claude-3.5-haiku, GPT-4o; Reasoning models: DeepSeek-R1, Gemini-2.5-pro, Claude-3.7-sonnet, GPT-o3. We also evaluated traditional static analysis tools, Slither and Mythril, on this Real-World Contracts Set. Consistent with findings in related work durieux2020empirical; sun2024gptscan, these tools exhibited very limited capability in detecting the complex logic vulnerabilities, often failing to identify any of the nuanced issues present.
@@ 4.4.6 RQ6: Comparison with Other LLM-based Auditing Methods
To contextualize SmartAuditFlow’s capabilities within the evolving landscape of LLM-assisted smart contract auditing, we compare its performance against several notable prior works. Notably, David et al. david2023you, PropertyGPT liu2024propertygpt, and GPTScan sun2024gptscan have demonstrated promising results using LLMs in different auditing scenarios.
David et al. david2023you employed direct prompting of GPT-4 and Claude-v1.3 for a binary classification task on 146 vulnerabilities across 38 types, identifying 58 TPs. PropertyGPT liu2024propertygpt focused on verifying LLM outputs, correctly identifying 9 out of 13 CVEs (focused on 5 classic types) and 17 out of 24 vulnerabilities from their SmartInv dataset. GPTScan sun2024gptscan targeted 10 common logic vulnerabilities, reporting 40 TPs. Their stated recall of 83.33% suggests these 40 TPs were out of approximately 48 actual instances of those types within their specific evaluation context. In contrast, SmartAuditFlow-Enhanced was evaluated on the 784 high and medium severity vulnerability instances within our challenging Real-World Contracts Set, identifying 310 TPs. A key distinction of our work is the aim to address a broad spectrum of vulnerability types ("All Types"), rather than a predefined subset of common or logic vulnerabilities.
To facilitate a more direct comparison on a common ground, we evaluated SmartAuditFlow-Enhanced on the set of 13 representative CVEs previously analyzed by PropertyGPT liu2024propertygpt. We compare our results with those reported for PropertyGPT, GPTScan, Slither, and Mythril (sourced from PropertyGPT’s study), and further benchmark against standalone invocations of reasoning-enhanced LLMs: GPT-o3 and DeepSeek-R1. The detailed detection results are presented in Table 5. SmartAuditFlow-Enhanced correctly detected all 13 out of 13 CVEs in this benchmark, achieving a 100% detection rate. This includes complex cases that other methods reportedly missed (e.g., CVE-2021-3004, CVE-2018-17111, as noted for PropertyGPT which detected 9/13).
@@ 5.1 Related Work
Recent studies have demonstrated growing potential for applying LLMs to software vulnerability analysis and smart contract security. Liu et al. liu2024exploring examined GPT’s performance across six tasks—including bug report summarization, security bug report identification, and vulnerability severity evaluation. They compared GPT’s results with 11 SOTA approaches. Their findings indicate that, when guided by prompt engineering, GPT can outperform many existing techniques. In parallel, Yin et al. yin2024multitask established benchmark performance metrics for open-source LLMs across multiple vulnerability analysis dimensions. They emphasized how strategic prompt design and model fine-tuning substantially impact detection accuracy and localization precision.
Researchers have extended these investigations to domain-specific languages (DSLs) likeC/C++, Java, and Solidity. Khare et al. khare2023understanding conducted an comprehensive evaluation of 16 advanced LLMs in the context of C/C++ and Java, showing that LLM-based approaches can surpass traditional static analysis and deep learning tools. Li et al. li2024llm integrated GPT-4 with CodeQL, a static analysis tool, to detect vulnerabilities in Java code, achieving an accuracy improvement of over 28% compared to using CodeQL alone.
Building on these foundations, recent advancements employ auxiliary techniques to enhance LLM capabilities. LLM4Vuln sun2024llm4vuln employs techniques such as knowledge retrieval, context supplementation, and advanced prompt schemes to boost vulnerability detection across multiple languages, including Java, C/C++, and Solidity. Their results show that GPT-4, when leveraging these enhancements, significantly outperforms competing models like Mixtral and CodeLlama. Similarly, GPTScan sun2024gptscan demonstrates that a binary (Yes/No) response format can be effectively used by GPT models to confirm potential vulnerabilities in Solidity smart contracts.
Furthermore, PropertyGPT liu2024propertygpt integrates GPT-4 with symbolic execution at the source code level, enabling formal verification of smart contract properties. Meanwhile, Ma et al. ma2024combining propose a paradigm-shifting architecture that decouples detection and reasoning tasks through two-stage fine-tuning. Simultaneously, they improved both accuracy and model interpretability in smart contract analysis.
@@ 4. Multi-Criteria Scoring Function $f(\rho,c,A)$:
For each component correctly identified (i.e., true positives from the $f_{coverage}$ step), assess the accuracy and completeness of its description in $\mathcal{M}(\rho,c)$ compared to the detailed description in $A$. Inspired by RAGAS es2024ragas, use another powerful LLM (the "judge model") to evaluate the quality of the primary LLM’s description against the expert analysis $A$. The judge model would be prompted with the primary LLM’s description of a component (e.g., function foo), the expert description of foo from $A$, and a rubric.

### v2

@@ 3.3.1 Execution and Reporting (A3-A5)
The final stage focuses on producing a clear, concise, and actionable audit report. This report consolidates the initial contract understanding ($s_{1}$), the audit plan ($s_{2}$), and the synthesized findings ($s_{4}$), including severity assessments and, where possible, actionable recommendations for mitigation. The structure and language of the report can be tailored to suit different audiences (e.g., developers requiring technical details, or management needing high-level summaries). Drawing insights from best practices in security reporting and relevant studies Liu et al. (2024); Yin et al. (2024), the report aims to cover key dimensions such as clear vulnerability descriptions, severity evaluations, precise localization of issues in the code, and concise summaries.
@@ 3.3.2 Task-Specific Review and Calibration
Given that LLMs can occasionally "hallucinate" or generate outputs that deviate from factual contract code Bender et al. (2021). Stage A3 is explicitly structured into two critical steps for each sub-task: (1) initial task-specific review and (2) rigorous calibration. This two-step process is designed to maximize accuracy and minimize false positives or missed vulnerabilities before findings are aggregated.
@@ 3.3.3 RAG-Enhanced Calibration for Improved Accuracy and Grounding
While pre-trained LLMs possess extensive general knowledge, their understanding of domain-specific knowledge can be limited or outdated. This can lead to gaps in analysis or less precise explanations, particularly for edge-case scenarios. To mitigate this, we enhance the calibration step with Retrieval-Augmented Generation (RAG) Lewis et al. (2020), as illustrated in Figure 2. RAG augments the LLM’s internal knowledge by providing access to relevant excerpts from external, up-to-date knowledge sources during generation.
We have observed instances where an LLM might correctly identify a vulnerable code location but provide an incomplete or imprecise explanation, especially for complex or uncommon vulnerabilities. RAG addresses this by retrieving documents that offer detailed, authoritative explanations for similar issues. Within our Execute Phase, each calibration prompt $\mathcal{P}_{v}^{i}$ is augmented with information retrieved from a curated knowledge base. This knowledge base includes smart contract databases (e.g., SWC Registry, CWE, CVE), security guidelines (e.g., Ethereum Foundation, ConsenSys Diligence), seminal research papers, technical blogs detailing novel exploits, and community-vetted best practice repositories (see Appendix B for a list of sources). These resources are embedded and stored in a Vector Database (using models like Sentence-BERT Reimers and Gurevych (2019)) for efficient semantic retrieval.
@@ 3.4 LLM-Powered Audit Evaluator
Assessing the quality of the generation is an even more arduous task than the generation itself, and this issue has not been given adequate consideration. Few studies have pointed that LLM has been a good evaluator for natural language generation (NLG), and the existing approaches are not tailored to the specific needs of the audit process Fu et al. (2023); Liu et al. (2023b). Inspired by RAGAs Es et al. (2024), we designed and implemented an automated LLM-Powered Audit Evaluator. This evaluator systematically compare the findings generated by SmartAuditFlow against expert-annotated ground truth answers from the same expert audited-report. The primary objectives of this evaluator are to ensure fair and reproducible assessments, accelerate the evaluation cycle, and provide actionable insights for the continual improvement of the SmartAuditFlow framework. Our LLM-Powered Audit Evaluator operates in two main phases: (1) LLM-Driven Finding Comparison and (2) Quantitative Performance Scoring.
@@
  • Match Categorization: Based on the multi-dimensional assessment and the rubric, the LLM categorizes the relationship between the paired findings. Inspired by prior research Wei et al. (2024), we adopt a nuanced classification approach rather than a binary true/false distinction. The primary categories include Exact Match (all critical aspects align), Partial Match (core issue identified but differs on a specific dimension like severity or location precision), and Incorrect (the response misidentifies the core issue in the Standard Answer).
@@ 4.2.1 Dataset
  • Standard Vulnerability Set: This dataset contains smart contracts annotated with common vulnerability types, which serves as a benchmark for evaluating the detection of well-understood vulnerability types against established tools. We utilize the SmartBugs-curated dataset Durieux et al. (2020), which is widely adopted by both developers and researchers. The dataset includes 143 contracts, each annotated with common vulnerabilities based on the DASP classification, making it an excellent resource for benchmarking vulnerability detection tools.
  • Real-world Contracts Set: To assess performance on more complex, real-world scenarios, we use a curated subset of contracts from projects audited by Code4rena Durieux et al. (2020). This set consists of 72 projects, encompassing 6,454 contracts, among which 243 contracts have known issues, including 784 instances of high or medium-severity vulnerabilities. This dataset tests SmartAuditFlow’s capability to handle complex contract logic and diverse vulnerability manifestations.
  • CVE Set: We curated a set of Common Vulnerabilities and Exposures (CVEs) relevant to smart contracts. As of early 2025, there are 592 smart contract-related CVEs, with integer overflows being a predominant category.Followed by another co-worker’s PropertyGPT Liu et al. (2025), we selected 13 representative CVEs for detailed analysis, including three integer overflow cases, three access control vulnerabilities, and a diverse set of four other logic bugs.
@@ 4.3 Implementation Details
All experiments for SmartAuditFlow and static tools were conducted within a Docker container to ensure a consistent, isolated, and reproducible environment across all test runs. For evaluating the vulnerability detection capabilities involving LLMs, we adopted a single attempt approach (often denoted as pass@1) Wei et al. (2024). This reflects practical auditing scenarios where auditors typically seek an effective assessment in one primary attempt.
@@ 4.4.3 RQ3: Performance on Real-World Smart Contracts
Based on findings from previous experiments (RQ1, RQ2) and their prevalence, we selected a representative set of standalone LLMs for comparison: General-Purpose models: DeepSeek-v3, Gemini-2.0-flash, Claude-3.5-haiku, GPT-4o; Reasoning models: DeepSeek-R1, Gemini-2.5-pro, Claude-3.7-sonnet, GPT-o3. We also evaluated traditional static analysis tools, Slither and Mythril, on this Real-World Contracts Set. Consistent with findings in related work Durieux et al. (2020); Sun et al. (2024a), these tools exhibited very limited capability in detecting the complex logic vulnerabilities, often failing to identify any of the nuanced issues present.
@@ 4.4.6 RQ6: Comparison with Other LLM-based Auditing Methods
To contextualize SmartAuditFlow’s capabilities within the evolving landscape of LLM-assisted smart contract auditing, we compare its performance against several notable prior works. Notably, David et al. David et al. (2023), PropertyGPT Liu et al. (2025), and GPTScan Sun et al. (2024a) have demonstrated promising results using LLMs in different auditing scenarios.
David et al. David et al. (2023) employed direct prompting of GPT-4 and Claude-v1.3 for a binary classification task on 146 vulnerabilities across 38 types, identifying 58 TPs. PropertyGPT Liu et al. (2025) focused on verifying LLM outputs, correctly identifying 9 out of 13 CVEs (focused on 5 classic types) and 17 out of 24 vulnerabilities from their SmartInv dataset. GPTScan Sun et al. (2024a) targeted 10 common logic vulnerabilities, reporting 40 TPs. Their stated recall of 83.33% suggests these 40 TPs were out of approximately 48 actual instances of those types within their specific evaluation context. In contrast, SmartAuditFlow-Enhanced was evaluated on the 784 high and medium severity vulnerability instances within our challenging Real-World Contracts Set, identifying 310 TPs. A key distinction of our work is the aim to address a broad spectrum of vulnerability types ("All Types"), rather than a predefined subset of common or logic vulnerabilities.
To facilitate a more direct comparison on a common ground, we evaluated SmartAuditFlow-Enhanced on the set of 13 representative CVEs previously analyzed by PropertyGPT Liu et al. (2025). We compare our results with those reported for PropertyGPT, GPTScan, Slither, and Mythril (sourced from PropertyGPT’s study), and further benchmark against standalone invocations of reasoning-enhanced LLMs: GPT-o3 and DeepSeek-R1. The detailed detection results are presented in Table 5. SmartAuditFlow-Enhanced correctly detected all 13 out of 13 CVEs in this benchmark, achieving a 100% detection rate. This includes complex cases that other methods reportedly missed (e.g., CVE-2021-3004, CVE-2018-17111, as noted for PropertyGPT which detected 9/13).
@@ 5.1 Related Work
Recent studies have demonstrated growing potential for applying LLMs to software vulnerability analysis and smart contract security. Liu et al. Liu et al. (2024) examined GPT’s performance across six tasks—including bug report summarization, security bug report identification, and vulnerability severity evaluation. They compared GPT’s results with 11 SOTA approaches. Their findings indicate that, when guided by prompt engineering, GPT can outperform many existing techniques. In parallel, Yin et al. Yin et al. (2024) established benchmark performance metrics for open-source LLMs across multiple vulnerability analysis dimensions. They emphasized how strategic prompt design and model fine-tuning substantially impact detection accuracy and localization precision.
Researchers have extended these investigations to domain-specific languages (DSLs) likeC/C++, Java, and Solidity. Khare et al. Khare et al. (2023) conducted an comprehensive evaluation of 16 advanced LLMs in the context of C/C++ and Java, showing that LLM-based approaches can surpass traditional static analysis and deep learning tools. Li et al. Li et al. (2024) integrated GPT-4 with CodeQL, a static analysis tool, to detect vulnerabilities in Java code, achieving an accuracy improvement of over 28% compared to using CodeQL alone.
Building on these foundations, recent advancements employ auxiliary techniques to enhance LLM capabilities. LLM4Vuln Sun et al. (2024b) employs techniques such as knowledge retrieval, context supplementation, and advanced prompt schemes to boost vulnerability detection across multiple languages, including Java, C/C++, and Solidity. Their results show that GPT-4, when leveraging these enhancements, significantly outperforms competing models like Mixtral and CodeLlama. Similarly, GPTScan Sun et al. (2024a) demonstrates that a binary (Yes/No) response format can be effectively used by GPT models to confirm potential vulnerabilities in Solidity smart contracts.
Furthermore, PropertyGPT Liu et al. (2025) integrates GPT-4 with symbolic execution at the source code level, enabling formal verification of smart contract properties. Meanwhile, Ma et al. Ma et al. (2024) propose a paradigm-shifting architecture that decouples detection and reasoning tasks through two-stage fine-tuning. Simultaneously, they improved both accuracy and model interpretability in smart contract analysis.
@@ 4. Multi-Criteria Scoring Function $f(\rho,c,A)$:
For each component correctly identified (i.e., true positives from the $f_{coverage}$ step), assess the accuracy and completeness of its description in $\mathcal{M}(\rho,c)$ compared to the detailed description in $A$. Inspired by RAGAS Es et al. [2024], use another powerful LLM (the "judge model") to evaluate the quality of the primary LLM’s description against the expert analysis $A$. The judge model would be prompted with the primary LLM’s description of a component (e.g., function foo), the expert description of foo from $A$, and a rubric.